Assume the Pacific Ocean has a surface area of 1.5 x 10^8 km^2 and an average depth of 3400 m. Calculate the volume of water in the ocean in liters.
Answer #1

As we already know, volume of a body can be calculated by the formula : Area of cross section * Thickness of cross section

Therefore, Volume of water in ocean = 1.5e8km2 * 3400 m = 1.5*108 * 106 m2 * 3400m [As 1km2 = 106m2, 1.5e8km2 is converted]

= 1.5*1014 * 3400 m3

= 5100*1014 m3 = 5.1*1011 m3

Since 1m3 = 1000L,

5.1*1011 m3 = 5.1*1011 * 1000 L

= 5.1*1014 L

Please let me know if you have any doubts. If not, please leave a positive rating :)